OpenStudy (anonymous):

If you were given:\[f(x) = \frac{1}{\frac{1}{x}}\] would the domain be (-∞, ∞) or (-∞, 0)U(0, ∞)? This is a question for all such examples where you must multiply by the reciprocal in some way.

OpenStudy (anonymous):

I think it would be the latter, \((-\infty,0)\cup(0,\infty)\). With the way the function is defined, \(x=0\) yields an indeterminate result. This despite \(\dfrac{1}{\frac{1}{x}}=x\) for \(x\not=0\). It's like with the following function: \[f(x)=\frac{x}{x}\] For nonzero \(x\), \(f(x)=1\). The plot of \(f(x)\) would be a horizontal line with a hole at \(x=0\). For your function, the plot would be that of \(f(x)=x\) with a hole at \(x=0\), I think.

OpenStudy (anonymous):

Ah. That makes sense. I was always wondered whether you simplified algebraically first or if you took the domain of the original. Same concept of removable discontinuities and such. Thank you :)

OpenStudy (anonymous):

For confirmation, WolframAlpha agrees: http://www.wolframalpha.com/input/?i=1%2F%281%2Fx%29 You're welcome.